What is 2 3/8 divided by 1/4
Answer by stanbon(75887)

What is 2 3/8 divided by 1/4
------
= [(8*2+3)/8] / [1/4]
--------
= (19/8) / (1/4)
--------
Invert the denominator and multiply:
------
= (19/8) * (4/1)
-----
= 19/2
=============
Cheers,
Stan H.
=============